OpenStudy (anonymous):

Let L be the line with parametric equations: x=5-2t y=-10+t z=-2+2t Find the vector equation for a line that passes through the point P=(2, 9, 9) and intersects L at a point that is distance 3 from the point Q=(5, −10, −2). Note that there are two possible correct answers.

OpenStudy (anonymous):

x= 2+t(?) y=9+t(?) z=9+t(?) This is all i know so far

OpenStudy (anonymous):

well you need to find those two points on the line that are 3 units away from point Q

OpenStudy (anonymous):



OpenStudy (anonymous):

essentially you have: 3^2 = (x - 5)^2 + (y + 10)^2 + (z + 2)^2 and you know what x,y,z are

OpenStudy (anonymous):

solve for t. Then use it find the two points on the line. Then find the vectors of the line and you'll have the answer

OpenStudy (anonymous):

Answers were: x=2+5t y=9-20t z=9-13t and x=2+t y=9-18t z=9-9t Still don't understand it though :(

OpenStudy (anonymous):

did you solve for t?

OpenStudy (anonymous):

3^2 = (5-2t - 5)^2 + (-10+t + 10)^2 + (-2+2t + 2)^2 ^ this is what I meant when I said solve for t

OpenStudy (anonymous):

https://www.wolframalpha.com/input/?i=3%5E2+%3D+%285-2t+-+5%29%5E2+%2B+%28-10%2Bt+%2B+10%29%5E2+%2B+%28-2%2B2t+%2B+2%29%5E2 so t = +/-1 when t = 1 x=5-2(1) = 3 y=-10+1 = -9 z=-2+2(1) = 0 when t = -1 x=5-2(-1) = 7 y=-10+(-1) = -11 z=-2+2(-1) = -4 so the two points that are 3 units away from Q are (3,-9,0) and (7,-11,-4)

OpenStudy (anonymous):

the two vectors are: (3-2,-9-9,0-9) = (1,-18,-9) and (7-2,-11-2,-4-2) = (5,-13,-6)

OpenStudy (anonymous):

oops typo (7-2,-11-9,-4-9) = (5,-20,-13)

OpenStudy (anonymous):

r1(t) = t(1,-18,-9) + (2,9,9) r2(t) = t(5,-20,-13) + (2,9,9) which match the solution

OpenStudy (anonymous):

Cool, makes sense. Thanks!